OpenStudy (anonymous):

Find all solutions of the equation 2\sin^2 x -\cos x =1 Apparently there are 3 solutions, but I don't know how to find them?

OpenStudy (thomas5267):

\[ \sin^2(x)+\cos^2(x)=1\\ \sin^2(x)=1-\cos^2(x) \]

OpenStudy (anonymous):

OOHHH ok thank you! can't believe I didn't think of that